OpenStudy (anonymous):

Multiply. Express your answer in lowest terms. 1/3 * 2/9 =

OpenStudy (anonymous):

@THEUSER822

OpenStudy (anonymous):

@Secret-Ninja

OpenStudy (anonymous):

Okay here is my work

OpenStudy (anonymous):

1*2 = 2

OpenStudy (secret-ninja):

\[\frac{ 1 }{ 3 } \times \frac{ 2 }{ 9 }\]

OpenStudy (anonymous):

3*9= 27

OpenStudy (secret-ninja):

That?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@TheAsker2002 Correct.

OpenStudy (anonymous):

So its 2/27

OpenStudy (anonymous):

Do you need the answer in decimal or fraction?

OpenStudy (secret-ninja):

You're right. :D Good job ^-^

OpenStudy (anonymous):

But... do you divide to make it smaller?

OpenStudy (anonymous):

its going to be in fractions @Linchpin

OpenStudy (anonymous):

Yes, divide it.

OpenStudy (secret-ninja):

You divide it to make it a decimal, and you simplify it to make it smaller.

OpenStudy (anonymous):

Can you use a calculator?

OpenStudy (secret-ninja):

You have to find out if you can divide the top and bottom numbers by the same number.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

could it be 3?

OpenStudy (anonymous):

2 can only be divisible by even numbers, So just divide it using your calculator.

OpenStudy (secret-ninja):

You can divide both by 1, but that will give you what you already have. If you divide by 2, 27 will be a decimal, so it won't work. 2/27 is already simplified.

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

So just use the calculator 2/ 27 =0.074

OpenStudy (secret-ninja):

^ thats for the decimal. :)

OpenStudy (anonymous):

Yeah, I was just telling her. She already has the answer in fraction i.e =22/7

OpenStudy (anonymous):

Thanks

OpenStudy (secret-ninja):

No problem. :)

OpenStudy (anonymous):

yw. ;)